# Simon Briend

### v2

@@ 2.3 Performance of Jordan ordering in the URRT model
Let $T\sim\text{{\sc urrt}}$. Then there exist positive constants $C$, $K$ such that for $21\leq\alpha<$ 2 $R_{\alpha}(\widehat{\sigma}_{J})\leq K(\alpha)n^{2-\alpha}+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{4}(n)~{},$ where $K(\alpha)=\frac{2}{2-\alpha}+\frac{2e^{2}}{(2-\alpha)^{2}}+\frac{2}{(2-\alpha)^{3}}$. Moreover, for $\alpha\geq 2$ $R_{\alpha}(\widehat{\sigma}_{J})\leq C\log^{4}(n)~{}.$
@@
Let $T\sim\text{{\sc urrt}}$. For $\alpha=1$ $R_{\alpha}(\widehat{\sigma}_{J})\leq\left(1+o(1)\right)1170R^{*}_{1}$ and for $\alpha\in(1,2)$, $R_{\alpha}(\widehat{\sigma}_{J})\leq\left(1+o(1)\right)\left(\frac{1}{2-\alpha}+\frac{3}{(2-\alpha)^{2}}+\frac{1}{(2-\alpha)^{3}}\right)70R_{\alpha}^{*}~{}.$
@@
$R_{\alpha}(\widehat{\sigma}^{\prime})\leq\sum_{i=1}^{n}\mathbb{E}\left[\frac{\widehat{\sigma}^{\prime}(i)}{i^{\alpha}}\right]+\sum_{i=1}^{n}\frac{i}{i^{\alpha}}~{}.$
@@
$\mathbb{E}\left[\widehat{\sigma}^{\prime}(i)\right]\leq\mathbb{E}\left[\sum_{j=1}^{n}\mathbbm{1}_{{\rm de}(j)\geq{\rm de}(i)}\right]~{}.$ (11)
@@
$\mathbb{P}\left\{{\rm de}(j)\geq{\rm de}(i)\right\}\leq\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{i,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau_{i,j}\right\}~{}.$
@@
$\mathbb{E}\left[\widehat{\sigma}^{\prime}(i)\right]\leq i+1+\sum_{j>i+1}\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{i,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau_{i,j}\right\}~{}.$ (12)
@@
$\mathbb{P}\left\{k\in(T,1)_{j}\right\}=\frac{de_{k-1}(j)+1}{k-1}~{},$
and each connection of a new vertex to a descendant of $j$ is independent of the previous ones, conditionally on ${\rm de}(j)$. Let $N_{n}:=n-j$ and let $\widetilde{W}_{N}={\rm de}(j)$ be the number of descendants of $j$ at time $n$. We thus have that $\widetilde{W}_{N}={\rm de}(j)$ follows a Pólya urn distribution, where the Pólya urn process has balls of two colours, it is started when $j$ is added to $T$, and it is run for a maximum number of steps $N_{n}$. From Mahmoud [22, Section 3.2], we have that
$\mathbb{E}\left[{\rm de}(j)+1\right]\ =\ \frac{n}{j}~{},$
@@
$\mathbb{P}\left\{{\rm de}(j)=k\right\}\ =\ \frac{k!(j-1)(j)\cdots(n-k-2)}{j(j+1)\cdots(n-1)}\binom{n-j}{k}~{}.$ (13)
@@
$\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{1,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(1)}{n}\leq\tau_{1,j}\right\}\leq\frac{1}{j}\left(2e^{2}+\log(j)\right)+\frac{1}{n-1}~{}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i}\right]\leq 18n~{}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq\left(\frac{2}{2-\alpha}+\frac{2e^{2}}{(2-\alpha)^{2}}+\frac{2}{(2-\alpha)^{3}}\right)n^{2-\alpha}~{}.$
@@ 3.1 A lower bound
In the pa model, we have, for $\alpha>0$ and $n\geq 300$ $R_{\alpha}^{*}\geq\frac{n^{2-\alpha}}{70}~{}.$
@@ 3.2 Performance of the Jordan ordering in the PA model
Let $T\sim\text{{\sc PA}}$. Then, there exist positive constants $C,\ K$, such that for $\alpha\in[1,3/2)$ $R_{\alpha}(\widehat{\sigma}_{J})\leq K(\alpha)n^{2-\alpha}+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{2}(n)\sqrt{n}~{},$ where $K(\alpha)=\frac{2}{2-\alpha}+\frac{8\sqrt{2}+\frac{10}{\sqrt{2}}}{(3/2-\alpha)(2-\alpha)}+\frac{20}{\sqrt{2}(2-\alpha)(3/2-\alpha)^{2}}$. For $\alpha\geq 3/2$, $R_{\alpha}(\widehat{\sigma}_{J})\leq cn^{3/2}~{},$ where $c$ is a positive constant.
Corollary 9 is a direct consequence of Theorems 7 and 8. It states that the Jordan ordering has a risk of optimal order for $\alpha\in[1,3/2)$. Let us remark that, here, the boundary value $3/2$ does not appear for the same reason as in the urrt case. In the urrt, the optimality result is limited to $\alpha<2$ because of the error originating from the estimation of vertex $1$. Here, the limitation to $\alpha<3/2$ has a different origin than in the URRT model. Indeed, the descendant ordering is only order optimal for $\alpha<3/2$, meaning that even if the position of vertex $1$ was known, ordering vertices by the number of their descendants would not result in a risk bound that matches the lower bound for $\alpha\geq 3/2$.
@@
Let $T\sim\text{{\sc PA}}$. For $\alpha\in[1,3/2)$ $R_{\alpha}(\widehat{\sigma}_{J})\leq\left(1+o(1)\right)70\left(\frac{2}{2-\alpha}+\frac{8\sqrt{2}+\frac{10}{\sqrt{2}}}{(3/2-\alpha)(2-\alpha)}+\frac{20}{\sqrt{2}(2-\alpha)(3/2-\alpha)^{2}}\right)R_{\alpha}^{*}~{}.$
@@ 4 Simulations
In this section, we first report a numerical illustration of our theoretical results on trees generated from the urrt and pa models. Then, we compare the performance of the Jordan ordering to other ordering procedures. For computational reasons, we display results for the descendant ordering procedure. The descendant ordering can be computed in time $\mathcal{O}(n\log(n))$. Also, one can find the Jordan centroid in linear time.
@@
In the second experiment, we perform an empirical comparison of Jordan ordering with the three following ordering methods:
@@
Similarly to the previous experiment, we compute the risk $R_{\alpha}$ for the four methods, on trees simulated from the urrt or pa models, in multiple settings. From Figures 5 and 6 we see that the Jordan estimator has the lowest risk, for all values of the trees sizes, and that the degree method is the second best one. In fact, it is not surprising that the degree method performs well for pa trees, since, in this model, the degree has a power law distribution and the order by degree correlates well with the arrival times of the vertices. However, this results is more surprising for the urrt model, where degree-centrality is known to be sub-optimal in the root-finding problem, see Bubeck et al. [6]. This is discussed further in Appendix A.3. Moreover, the spectral method has the poorest performance in both models. A possible explanation for this is that the random fluctuations of the adjacency matrix in the considered recursive tree models are large, leading to a large difference between the expected and empirical adjacency matrices in spectral norm. This absence of concentration, which is generally required in spectral ordering methods, could explain why this method poorly performs in our setting. Finally, the Reverse DMC algorithm performs similarly poorly as the spectral method in the pa model.
@@ A.1 A remark on the choice of the risk
$R_{\alpha}(\Sigma)\leq c_{\alpha}n^{2-\alpha}~{},$
for some positive constant $c_{\alpha}$. On the other hand, Theorems 1 and 7 imply that for $\alpha<1$, $R_{\alpha}^{*}\geq c^{\prime}_{\alpha}n^{2-\alpha}$. Therefore, with $\alpha<1$, the random ordering has a risk of the same order as that of the optimal one. This is why we restrict our analysis of the risk to $\alpha\geq 1$. Our analysis of the Jordan ordering proves that this method has a risk of optimal order for $\alpha\in[1,2)$ in the urrt case and $\alpha\in[1,3/2)$ in the pa tree.
It is also possible to change the risk by replacing $1/\sigma(i)^{\alpha}$ by $f(\sigma(i))$, for some non-increasing function $f$. An example is $f(t)=\mathbb{I}_{\{t\leq\tau_{n}\}}$, for some threshold $\tau_{n}$. Such a risk measures the quality of the ordering on the first $\tau_{n}$ vertices only. For many choices of $f$, our analysis can be directly adapted However, for rapidly decreasing functions (such as the step function above), the Jordan ordering is not be optimal up to a constant factor. This suggests that one should investigate methods that are better at ordering the early vertices.
@@ A.3 A remark on ordering by degree
As discussed in Section 4, a simple ordering procedure is by the degrees. Simulations suggest that it does not perform as well as Jordan ordering, and it may produces non-recursive ordering. Nonetheless, it is a simple procedure worth mentioning. Since in a pa tree the degree of a given node follows a Pólya urn distribution, analysing the performance of the degree ordering is similar to the analysis carried out for Jordan centrality. However, the simulations results displayed in Figure 7 suggest that for any $\alpha\in[1,3/2)$ the risk of the degree ordering grows at a faster rate than $n^{2-\alpha}$. Both in the URRT and PA model, for sizes of trees $\{1000,2000,4000,8000\}$, we sample $10$ trees at each size and compute the risk of the descendant and degree ordering for different values of $\alpha$. Then, for each value of $\alpha$, we perform a linear regression on the log-plot of the risk, to estimate the exponent of the polynomial.
@@
On the other hand, ordering vertices by their degree in the urrt is known to be suboptimal for finding the root, as there are many vertices with much higher degree (Eslava [13]). Simulation results displayed in Figure 8 suggest that, for most values of $\alpha$, the risk of ordering by degree in the urrt model grows at a faster rate than $n^{2-\alpha}$ for any $\alpha\in(1,2)$. On the other hand, observing Figure 8, it seems like, for $\alpha=1$, the degree ordering may have a risk growing at the optimal rate of $n$.
Figure 8: Estimation of the rate at which the risk increases with the size of the tree for different values of $\alpha$ in the urrt. Here we compare the case of descendant and degree ordering. Proposition 1 shows that the optimal risk grows as $n^{2-\alpha}$, and that the Jordan and descendant ordering’s risks grow at the same rate (see Proposition 4). This experiment is in accordance with these results. For the degree ordering, this plot suggests that the risk grows at a rate faster than $n^{2-\alpha}$.
In Section 4, where the empirical performance of different ordering procedures are compared, the degree ordering is the method with second best performance. The above simulations suggest that all the other tested methods have risks growing at a faster rate than $n^{2-\alpha}$.
@@ A.4 Proof of Theorem 4
$\mathbb{P}\left\{{\rm de}(j)=k\right\}\ =\ \frac{k!(j-1)(j)\cdots(n-k-2)}{j(j+1)\cdots(n-1)}\binom{n-j}{k}~{}.$
@@
$\leq\ \frac{j-1}{n-1}\left(1-\frac{k}{n-2}\right)^{j-2}~{}.$
@@
$\leq\tau_{i,j}i\left(1-\frac{1}{n-2}\right)^{i-2}\leq\tau_{i,j}i~{}.$
@@
$\mathbb{P}\left\{\frac{{\rm de}(1)}{n}\leq\tau_{i,j}\right\}\leq\sum_{k\in[1,\tau_{i,j}n]}\frac{1}{n-1}\leq\tau_{i,j}+\frac{1}{n-1}~{}.$
@@
$\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{i,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau_{i,j}\right\}\leq 2e^{2}e^{-\log\frac{j}{i}}+\frac{i}{j}\log\frac{j}{i}\leq\frac{i}{j}\left(2e^{2}+\log\frac{j}{i}\right)~{},$
@@
$\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{1,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(1)}{n}\leq\tau_{1,j}\right\}\leq\frac{1}{j}\left(2e^{2}+\log(j)\right)+\frac{1}{n-1}~{}.$
@@
$\leq 2n+3+\log(n)+\sum_{j=3}^{n}\frac{1}{j}\sum_{i=1}^{j-2}\left(2e^{2}+\log\frac{j}{i}\right)~{}.$ (14)
@@
$\sum_{i=1}^{j-2}\log\frac{j}{i}\leq\log\left(\frac{j^{j}}{j!}\right)~{},$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i}\right]\leq 3+\log(n)+(2+2e^{2}+\log 3)n~{},$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i}\right]\leq 18n~{}.$
@@
$\leq\frac{2}{2-\alpha}n^{2-\alpha}+1+\zeta(\alpha)+\frac{2e^{2}}{(2-\alpha)^{2}}n^{2-\alpha}+\sum_{j=3}^{n}\frac{1}{j}\sum_{i=1}^{j-2}\frac{1}{i^{\alpha-1}}\log\frac{j}{i}~{}.$
@@
$\sum_{i=1}^{j-2}\frac{1}{i^{\alpha-1}}\log\frac{j}{i}\leq 2\int_{1}^{j}\frac{1}{t^{\alpha-1}}\log\left(\frac{j}{t}\right)~{},$
@@
$\sum_{i=1}^{j-2}\frac{1}{i^{\alpha-1}}\log\frac{j}{i}\leq\frac{2}{(2-\alpha)^{2}}j^{2-\alpha}~{}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq\left(\frac{2}{2-\alpha}+\frac{2e^{2}}{(2-\alpha)^{2}}+\frac{2}{(2-\alpha)^{3}}\right)n^{2-\alpha}~{}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq C~{},$ (15)

### v3

@@ 2.3 Performance of Jordan ordering in the URRT model
Let $T\sim\text{{\sc urrt}}$. Then there exist positive constants $C$, $K$ such that for $21\leq\alpha<$ 2
$R_{\alpha}(\widehat{\sigma}_{J})\leq K(\alpha)n^{2-\alpha}+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{4}(n),$
where $K(\alpha)=\frac{2}{2-\alpha}+\frac{2e^{2}}{(2-\alpha)^{2}}+\frac{2}{(2-\alpha)^{3}}$. Moreover, for $\alpha\geq 2$
$R_{\alpha}(\widehat{\sigma}_{J})\leq C\log^{4}(n).$
@@
Let $T\sim\text{{\sc urrt}}$. For $\alpha=1$
$R_{\alpha}(\widehat{\sigma}_{J})\leq\left(1+o(1)\right)1170R^{*}_{1}$
and for $\alpha\in(1,2)$,
$R_{\alpha}(\widehat{\sigma}_{J})\leq\left(1+o(1)\right)\left(\frac{1}{2-\alpha}+\frac{3}{(2-\alpha)^{2}}+\frac{1}{(2-\alpha)^{3}}\right)70R_{\alpha}^{*}.$
@@
$R_{\alpha}(\widehat{\sigma}^{\prime})\leq\sum_{i=1}^{n}\mathbb{E}\left[\frac{\widehat{\sigma}^{\prime}(i)}{i^{\alpha}}\right]+\sum_{i=1}^{n}\frac{i}{i^{\alpha}}.$
@@
$\mathbb{E}\left[\widehat{\sigma}^{\prime}(i)\right]\leq\mathbb{E}\left[\sum_{j=1}^{n}\mathbbm{1}_{{\rm de}(j)\geq{\rm de}(i)}\right].$ (11)
@@
$\mathbb{P}\left\{{\rm de}(j)\geq{\rm de}(i)\right\}\leq\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{i,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau_{i,j}\right\}.$
@@
$\mathbb{E}\left[\widehat{\sigma}^{\prime}(i)\right]\leq i+1+\sum_{j>i+1}\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{i,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau_{i,j}\right\}.$ (12)
@@
$\mathbb{P}\left\{k\in(T,1)_{j}\right\}=\frac{de_{k-1}(j)+1}{k-1},$
and each connection of a new vertex to a descendant of $j$ is independent of the previous ones, conditionally on ${\rm de}(j)$. Let $N_{n}:=n-j$ and let $\widetilde{W}_{N}={\rm de}(j)$ be the number of descendants of $j$ at time $n$. We thus have that $\widetilde{W}_{N}={\rm de}(j)$ follows a Pólya urn distribution, where the Pólya urn process has balls of two colours, it is started when $j$ is added to $T$, and it is run for a maximum number of steps $N_{n}$. From Mahmoud [23, Section 3.2], we have that
$\mathbb{E}\left[{\rm de}(j)+1\right]\ =\ \frac{n}{j},$
@@
$\mathbb{P}\left\{{\rm de}(j)=k\right\}\ =\ \frac{k!(j-1)(j)\cdots(n-k-2)}{j(j+1)\cdots(n-1)}\binom{n-j}{k}.$ (13)
@@
$\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{1,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(1)}{n}\leq\tau_{1,j}\right\}\leq\frac{1}{j}\left(2e^{2}+\log(j)\right)+\frac{1}{n-1}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i}\right]\leq 18n.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq\left(\frac{2}{2-\alpha}+\frac{2e^{2}}{(2-\alpha)^{2}}+\frac{2}{(2-\alpha)^{3}}\right)n^{2-\alpha}.$
@@ 3.1 A lower bound
In the pa model, we have, for $\alpha=1$ and $n\geq 300$ $R_{\alpha}^{*}\geq\frac{n^{2-\alpha}}{70}~{}.$
@@
4 Preferential attachment tree
4.1 A lower bound
Theorem 8.
In the pa model, we have, for $\alpha>0$ and $n\geq 300$
$R_{\alpha}^{*}\geq\frac{n^{2-\alpha}}{70}.$
The proof is deferred to Appendix A.5.
Remark.  As in the case of the urrt model, we have
$R_{\alpha}^{*}\geq\frac{1}{2},$
which improves on the result of Theorem 8 for $\alpha>2$.
@@ 4.2 Performance of the Jordan ordering in the PA model
Let $T\sim\text{{\sc PA}}$. Then, there exist positive constants $C,\ K$, such that for $\alpha\in[1,3/2)$
$R_{\alpha}(\widehat{\sigma}_{J})\leq K(\alpha)n^{2-\alpha}+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{2}(n)\sqrt{n},$
where $K(\alpha)=\frac{2}{2-\alpha}+\frac{8\sqrt{2}+\frac{10}{\sqrt{2}}}{(3/2-\alpha)(2-\alpha)}+\frac{20}{\sqrt{2}(2-\alpha)(3/2-\alpha)^{2}}$. For $\alpha\geq 3/2$,
$R_{\alpha}(\widehat{\sigma}_{J})\leq cn^{3/2},$
where $c$ is a positive constant.
Corollary 10 is a direct consequence of Theorems 8 and 9. It states that the Jordan ordering has a risk of optimal order for $\alpha\in[1,3/2)$. Let us remark that, here, the boundary value $3/2$ does not appear for the same reason as in the urrt case. In the urrt, the optimality result is limited to $\alpha<2$ because of the error originating from the estimation of vertex $1$. Here, the limitation to $\alpha<3/2$ has a different origin than in the URRT model. Indeed, the descendant ordering is only order optimal for $\alpha<3/2$, meaning that even if the position of vertex $1$ was known, ordering vertices by the number of their descendants would not result in a risk bound that matches the lower bound for $\alpha\geq 3/2$.
@@
Let $T\sim\text{{\sc PA}}$. For $\alpha\in[1,3/2)$
$R_{\alpha}(\widehat{\sigma}_{J})\leq\left(1+o(1)\right)70\left(\frac{2}{2-\alpha}+\frac{8\sqrt{2}+\frac{10}{\sqrt{2}}}{(3/2-\alpha)(2-\alpha)}+\frac{20}{\sqrt{2}(2-\alpha)(3/2-\alpha)^{2}}\right)R_{\alpha}^{*}.$
@@ 5 Simulations
In this section, we first report a numerical illustration of our theoretical results on trees generated from the urrt and pa models. Then, we compare the performance of the descendant ordering to other ordering procedures. For computational reasons, we display results for the descendant ordering procedure. The descendant ordering can be computed in time $\mathcal{O}(n\log(n))$. Also, one can find the Jordan centroid in linear time.
@@
In the second experiment, we perform an empirical comparison of descendant ordering with the three following ordering methods:
@@
Similarly to the previous experiment, we compute the risk $R_{\alpha}$ for the four methods, on trees simulated from the urrt or pa models, in multiple settings. From Figures 5 and 6 we see that the descendant estimator has the lowest risk, for all values of the trees sizes, and that the degree method is the second best one. In fact, it is not surprising that the degree method performs well for pa trees, since, in this model, the degree has a power law distribution and the order by degree correlates well with the arrival times of the vertices. However, this results is more surprising for the urrt model, where degree-centrality is known to be sub-optimal in the root-finding problem, see Bubeck et al. [6]. This is discussed further in Appendix C. Moreover, the spectral method has the poorest performance in both models. A possible explanation for this is that the random fluctuations of the adjacency matrix in the considered recursive tree models are large, leading to a large difference between the expected and empirical adjacency matrices in spectral norm. This absence of concentration, which is generally required in spectral ordering methods, could explain why this method poorly performs in our setting. Finally, the Reverse DMC algorithm performs similarly poorly as the spectral method in the pa model.
@@ A.1 A remark on the choice of the risk
$R_{\alpha}(\Sigma)\leq c_{\alpha}n^{2-\alpha},$
for some positive constant $c_{\alpha}$. On the other hand, Theorems 1 and 8 imply that for $\alpha<1$, $R_{\alpha}^{*}\geq c^{\prime}_{\alpha}n^{2-\alpha}$. Therefore, with $\alpha<1$, the random ordering has a risk of the same order as that of the optimal one. This is why we restrict our analysis of the risk to $\alpha\geq 1$. Our analysis of the Jordan ordering proves that this method has a risk of optimal order for $\alpha\in[1,2)$ in the urrt case and $\alpha\in[1,3/2)$ in the pa tree.
It is also possible to change the risk by replacing $1/\sigma(i)^{\alpha}$ by $f(\sigma(i))$, for some non-increasing function $f$. An example is $f(t)=\mathbb{I}_{\{t\leq\tau_{n}\}}$, for some threshold $\tau_{n}$. Such a risk measures the quality of the ordering on the first $\tau_{n}$ vertices only. For many choices of $f$, our analysis can be directly adapted However, for rapidly decreasing functions (such as the step function above), the Jordan ordering is not optimal up to a constant factor. This suggests that one should investigate methods that are better at ordering the early vertices.
@@ A.3 A remark on ordering by degree
As discussed in the above simulations, a simple ordering procedure is by the degrees. Simulations suggest that it does not perform as well as Jordan ordering, and it may produces non-recursive ordering. Nonetheless, it is a simple procedure worth mentioning. Since in a pa tree the degree of a given node follows a Pólya urn distribution, analysing the performance of the degree ordering is similar to the analysis carried out for the Jordan ordering. However, the simulations results displayed in Figure 7 suggest that for any $\alpha\in[1,3/2)$ the risk of the degree ordering grows at a faster rate than $n^{2-\alpha}$. Both in the URRT and PA model, for sizes of trees $\{1000,2000,4000,8000\}$, we sample $10$ trees at each size and compute the risk of the descendant and degree ordering for different values of $\alpha$. Then, for each value of $\alpha$, we perform a linear regression on the log-plot of the risk, to estimate the exponent of the polynomial.
@@
On the other hand, ordering vertices by their degree in the urrt is known to be suboptimal for finding the root, as there are many vertices with much higher degree ([14]). Simulation results displayed in Figure 8 suggest that, for most values of $\alpha$, the risk of ordering by degree in the urrt model grows at a faster rate than $n^{2-\alpha}$ for any $\alpha\in(1,2)$. On the other hand, observing Figure 8, it seems like, for $\alpha=1$, the degree ordering may have a risk growing at the optimal rate of $n$.
Figure 8: Estimation of the rate at which the risk increases with the size of the tree for different values of $\alpha$ in the urrt. Here we compare the case of descendant and degree ordering. Proposition 1 shows that the optimal risk grows as $n^{2-\alpha}$, and that the Jordan and descendant orderings’ risks grow at the same rate (see Proposition 4). This experiment is in accordance with these results. For the degree ordering, this plot suggests that the risk grows at a rate faster than $n^{2-\alpha}$.
In the Supplementary Material, where the empirical performance of different ordering procedures are compared, the degree ordering is the method with second best performance. The above simulations suggest that all the other tested methods have risks growing at a faster rate than $n^{2-\alpha}$.
@@ A.4 Proof of Theorem 4
$\mathbb{P}\left\{{\rm de}(j)=k\right\}\ =\ \frac{k!(j-1)(j)\cdots(n-k-2)}{j(j+1)\cdots(n-1)}\binom{n-j}{k}.$
@@
$\leq\ \frac{j-1}{n-1}\left(1-\frac{k}{n-2}\right)^{j-2}.$
@@
$\leq\tau_{i,j}i\left(1-\frac{1}{n-2}\right)^{i-2}\leq\tau_{i,j}i.$
@@
$\mathbb{P}\left\{\frac{{\rm de}(1)}{n}\leq\tau_{i,j}\right\}\leq\sum_{k\in[1,\tau_{i,j}n]}\frac{1}{n-1}\leq\tau_{i,j}+\frac{1}{n-1}.$
@@
$\mathbb{P}\left\{\frac{{\rm de}(j)}{n}\geq\tau_{1,j}\right\}+\mathbb{P}\left\{\frac{{\rm de}(1)}{n}\leq\tau_{1,j}\right\}\leq\frac{1}{j}\left(2e^{2}+\log(j)\right)+\frac{1}{n-1}.$
@@
$\leq 2n+3+\log(n)+\sum_{j=3}^{n}\frac{1}{j}\sum_{i=1}^{j-2}\left(2e^{2}+\log\frac{j}{i}\right).$ (14)
@@
$\sum_{i=1}^{j-2}\log\frac{j}{i}\leq\log\left(\frac{j^{j}}{j!}\right),$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i}\right]\leq 3+\log(n)+(2+2e^{2}+\log 3)n,$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i}\right]\leq 18n.$
@@
$\leq\frac{2}{2-\alpha}n^{2-\alpha}+1+\zeta(\alpha)+\frac{2e^{2}}{(2-\alpha)^{2}}n^{2-\alpha}+\sum_{j=3}^{n}\frac{1}{j}\sum_{i=1}^{j-2}\frac{1}{i^{\alpha-1}}\log\frac{j}{i}.$
@@
$\sum_{i=1}^{j-2}\frac{1}{i^{\alpha-1}}\log\frac{j}{i}\leq 2\int_{1}^{j}\frac{1}{t^{\alpha-1}}\log\left(\frac{j}{t}\right),$
@@
$\sum_{i=1}^{j-2}\frac{1}{i^{\alpha-1}}\log\frac{j}{i}\leq\frac{2}{(2-\alpha)^{2}}j^{2-\alpha}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq\left(\frac{2}{2-\alpha}+\frac{2e^{2}}{(2-\alpha)^{2}}+\frac{2}{(2-\alpha)^{3}}\right)n^{2-\alpha}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq C,$ (15)